# Lower Semicontinuity of Restricted Holonomy Groups

### v1

@@ 3. Application to Riemannian holonomy groups
The most useful use case of Theorem 2.4 is probably the case of oriented Riemannian manifolds and their Levi-Civita connection.
@@
Let $M$ be a manifold of class $C^{1}$.
  1. (1)
    We say that a map $f:M\to X$ into a normed vector space $X$ or a section $f:M\to X$ of a vector bundle with norm over $M$ is locally Lipschitz-$C^{1}$, if $f$ is of class $C^{1}$ and $df:TM\to TX$ is locally Lipschitz continuous.
  2. (2)
    Let $\mathrm{Met}(M)$ be the set of all locally Lipschitz-$C^{1}$-Riemannian metrics on $M$, equipped with the compact-open $C^{1}$-topology.
  3. (3)
    Given a Riemannian metric $h\in\mathrm{Met}(M)$ with Levi-Civita connection $\nabla^{h}$ and $x\in M$, we denote $\mathrm{Hol}_{x}(h):=\mathrm{Hol}_{x}(\nabla^{h})$ and $\mathrm{Hol}_{x}^{0}(h):=\mathrm{Hol}_{x}^{0}(\nabla^{h})$.
Now assume that we have a sequence of locally Lipschitz-$C^{1}$ Riemannian metrics converging in $C^{1}$. Then the condition of convergence $\nabla^{k}\to\nabla^{g}$ is automatically satisfied, as well as the condition on Lipschitz continuity of the connections, and we get the following corollary.
Corollary 3.2.
Let $(M,g)$ be an oriented $n$-dimensional Riemannian manifold of class $C^{1}$, let $g\in\mathrm{Met}(M)$ and for $k\in\mathbb{N}$, let $g_{k}\in\mathrm{Met}(M)$. Let $H\subseteq\mathrm{SO}(n)$ be a closed subgroup and let $x\in M$. Assume that $g_{k}\to g$ in $\mathrm{Met}(M)$ and for all $k\in\mathbb{N}$, we have $\mathrm{Hol}_{x}(g_{k})\subseteq H$ up to conjugation in $\mathrm{SO}(n)$. Then, $\mathrm{Hol}_{x}(g)\subseteq H$ up to conjugation in $\mathrm{SO}(n)$.
@@
Let $M$ be an oriented $n$-dimensional Riemannian manifold of class $C^{1}$, let $H\subseteq\mathrm{SO}(n)$ be a closed subgroup and let $x\in M$. Then, the set
@@
A useful consequence of Theorem 2.4 (respectively, the variant for the restricted holonomy groups) is the observation that the map assigning to a Riemannian metric the conjugacy class of the restricted holonomy group of its Levi-Civita connection is lower semicontinuous with respect to the $C^{1}$-topology and the inclusion order relation. The proof will follow as a straightforward application of Theorem 2.4, once it has been shown that this statement is actually well-defined.
@@
The relation $\leq$” on the set $\mathrm{CSC}(G)$ defined as above (for $G$ any compact Lie group) will be shown to define a partial order on $\mathrm{CSC}(G)$. Hence it is then well-defined to speak of lower semi-continuous functions to $\mathrm{CSC}(G)$. To prove antisymmetry, the following Cantor-Bernstein-like theorem is needed, whose proof has its origins in the one from the article [6] by Müller.
@@
At first, we prove that the linear map $d_{1}f:T_{1}K\to T_{1}H$ is injective. This can be seen as follows: It is well known that for the exponential maps $\exp_{K}$ and $\exp_{H}$ of $K$ resp. $H$ we have $\exp_{H}\circ d_{1}f=f\circ\exp_{K}$. Furthermore, both exponential maps are local diffeomorphisms, in particular injective if restricted to a suitable neighbourhood of zero in the respective tangent spaces.
@@
and therefore $w=0$ by injectivity of $(f\circ\exp_{K})|_{U}$. By linearity of $d_{1}f$, also $v=0$.
Thus $d_{1}f$ is injective, and analogously one shows that $d_{1}g:T_{1}H\to T_{1}K$ is also injective. By the Cantor-Bernstein theorem in the category of finite-dimensional vector spaces, both tangent spaces have the same dimension and both linear maps are isomorphisms.
Now let $H_{0},K_{0}$ be the connected component of $1\in H_{0}$ and $1\in K_{0}$, respectively. It is well known that $\exp_{H}$ surjects onto $H_{0}$; since furthermore $d_{1}f$ is surjective by the previous paragraph and by the formula $\exp_{H}\circ d_{1}f=f\circ\exp_{K}$ again, one concludes that $H_{0}\subseteq f(K_{0})$. Thus $f|_{K_{0}}:K_{0}\to H_{0}$ is an isomorphism. The map $f$ induces a map $\overline{f}:K/K_{0}\to H/H_{0}$ which fits into the following commutative diagram with exact rows:
@@
Thus for every $h\in H$, there exist $k_{1}\in K$ and $h_{0}\in H_{0}$ such that $h=f(k_{1})h_{0}$. On the other hand, due to surjectivity of $f|_{K_{0}}:K_{0}\to H_{0}$, there exists $k_{0}\in K_{0}$ with $h_{0}=f(k_{0})$, and then $h=f(k_{1})f(k_{0})=f(k_{1}k_{0})$ lies in the image of $f$. One concludes that $f$ is surjective, thus bijective.
Now to finish the proof, it suffices to note that $f^{-1}$ is also a Lie group homomorphism, which is well known and follows from the fact that $f$ induces an isomorphism of the Lie algebras as well and an application of the implicit function theorem.
@@
Now the announced reformulation of Theorem 2.4 for restricted Riemannian holonomy groups can be stated.
Theorem 3.7.
Let $M$ be an oriented $n$-dimensional manifold of class $C^{1}$ and let $x\in M$. The map
is lower-semicontinuous.
Proof.
It is well known (cf. e.g. [5], Theorem 3.2.8) that for any $h\in\mathrm{Met}(M)$, the group $\mathrm{Hol}_{x}^{0}(h)$ is closed in $\mathrm{SO}(n)$. Hence the map from the statement of the theorem is actually well-defined. Now the statement follows from Theorem 2.4. ∎
@@
For any $k\in\mathbb{N}$, $(M,g_{k})$ is up to scaling describable as a subset of the hyperbolic space of dimension $2$, which is the symmetric space $\mathrm{SO}_{0}(1,2)/(\{1\}\times\mathrm{SO}(2))$, where $\{1\}\times\mathrm{SO}(2)$ is the isotropy subgroup (cf. [3], Chapter 10.K, Table 3). Hence, as is well known, $\mathrm{Hol}_{x}(g_{k})=\mathrm{Hol}_{x}^{0}(g_{k})=\{1\}\times\mathrm{SO}(2)\cong\mathrm{SO}(2)$ for all $x\in M$ (cf. [3], Proposition 10.79). However $\mathrm{Hol}_{x}^{0}(g)=\{I_{2}\}\subset\mathrm{SO}(2)$ because $g$ is flat.
It is also easy to perturb a Calabi-Yau metric without leaving the realm of Kähler metrics, that is, obtain a sequence of metrics with holonomy $\mathrm{U}(n)$ converging to a metric with holonomy $\mathrm{SU}(n)$:
@@
Let $M$ be the Fermat quartic, which is a $\mathrm{K}3$ surface in $\mathbb{C}\mathrm{P}^{3}$. Let $\omega$ be the Kähler form of some Kähler metric on $M$. Let $\alpha\in\ker(\partial)\backslash\ker(\overline{\partial})\subset C^{\infty}(M,\Lambda^{1,0}T^{*}M)$. [^6] For example, let $h\in C^{\infty}(M,\mathbb{R})$ be such that $\overline{\partial}\partial h$ is not identically zero; such a function can be defined in a chart. Then, define $\alpha:=\partial h$. Let $\rho:=\overline{\partial}\alpha=d\alpha\in C^{\infty}(M,\Lambda^{1,1}T^{*}M)$; by construction $\rho$ represents zero in cohomology.
By the Calabi conjecture (solved by Yau in [11], [12]), there exists a unique Kähler metric $g_{1}$ on $M$ with Ricci form $\rho$ and such that its Kähler form is contained in the cohomology class of $\omega$. Moreover, there exists a unique Ricci-flat Kähler metric $g_{\infty}$ on $M$ with Kähler form contained in the cohomology class of $\omega$. Let $g_{k}:=\frac{1}{k}g_{1}+\frac{k-1}{k}g_{\infty}$ for $k\in\mathbb{N}_{\geq 2}$. Then, the metrics $g_{k}$ are Kähler since convex combinations of Kähler metrics are again Kähler, and the sequence $(g_{k})_{k\in\mathbb{N}}$ converges in $\mathrm{Met}(M)$ to $g_{\infty}$. Moreover, for all $k\in\mathbb{N}$ the metric $g_{k}$ is not Ricci-flat. [^7] If it were, it would contradict the uniqueness statement from the Calabi conjecture: The metrics $g_{k}$ also all have Kähler forms contained in the cohomology class of $\omega$. However, there exists only a unique such one which is Ricci-flat, namely $g_{\infty}$.
Thus, the metrics $g_{k}$ all have holonomy group contained in $\mathrm{U}(2)$ (by e. g. [5], proposition 4.4.2), but not in $\mathrm{SU}(2)$ (see e. g. [5], proposition 6.1.1). However, $g_{\infty}$ is Ricci-flat and hence has its holonomy group contained in $\mathrm{SU}(2)\subset\mathrm{U}(2)$.
@@
On the other hand, it would be impossible to construct an example as the previous one on a closed manifold: On a compact manifold $M$ of dimension $n$, all Einstein metrics are stationary points of the Einstein-Hilbert functional $S:\mathrm{Met}(M)\to\mathbb{R}$, $g\mapsto\frac{\int_{M}\mathrm{scal}^{g}\,\mathrm{dvol}^{g}}{\left(\int_{M}\mathrm{dvol}^{g}\right)^{(n-2)/n}}$ (where $\mathrm{scal}^{g}$ is the scalar curvature of the metric $g$ and $\mathrm{dvol}^{g}$ is its volume form). In particular, one sees from this that it is impossible to deform Ricci-flat Einstein metrics (such as hyperkähler ones) on a compact manifold to non-Ricci-flat Einstein metrics (such as quaternion-Kähler ones), since $S$ needs to remain constant along any deformation which does not leave the space of Einstein metrics. Contrarily, Theorem 3.7 holds without any compactness assumption on the manifold.
To conclude this paper, it shall be remarked that one would assume that the conditions on Theorem 2.4 can actually be weakened. The most apparent restriction is the one to connections compatible with positive definite metrics, which makes the theorem, e.g., inapplicable in the Lorentzian setting. Moreover, it would be interesting to see whether Theorem 3.7 also holds for the full holonomy group (which is, as is known due to an example made by Wilking in [10], in general not closed in the special orthogonal group).
However, the relaxation of these conditions would require entirely different proofs. It would be interesting to study such generalizations.

### v2

@@ 3. Application to Riemannian holonomy groups
Theorem 1.1 may be applied to and refined in the case of oriented Riemannian manifolds and their Levi-Civita connections.
@@
Let $M$ be a manifold of class $C^{2}$. Let $\mathrm{Met}(M)$ be the set of all $C^{1}$-Riemannian metrics on $M$, equipped with the compact-open $C^{1}$-topology. Given $h\in\mathrm{Met}(M)$ with Levi-Civita connection $\nabla^{h}$ and $x\in M$, we denote $\mathrm{Hol}_{x}(h):=\mathrm{Hol}_{x}(\nabla^{h})$ and $\mathrm{Hol}_{x}^{0}(h):=\mathrm{Hol}_{x}^{0}(\nabla^{h})$.
The proof of Corollary 1.2 is straightforward.
Proof of Corollary 1.2.
By the Koszul formula, $C^{1}$-convergence $g_{k}\to g$ implies $C^{0}$-convergence $\nabla^{g_{k}}\to\nabla^{g}$. The corollary follows from Theorem 1.1. ∎
@@
Let $M$ be an oriented $n$-dimensional Riemannian manifold of class $C^{2}$, let $H\subseteq\mathrm{SO}(n)$ be a closed subgroup and let $x\in M$. Then, the set
@@
For the restricted holonomy group, the results can also be phrased in terms of lower semicontinuity of the map $\mathrm{Met}(M)\to\mathrm{CSC}(\mathrm{SO}(n))$, $g\mapsto[\mathrm{Hol}_{x}^{0}(g)]$ as stated before, see Theorem 1.3. The proof will follow again as an application of Theorem 1.1, once it has been shown that the statement is actually well-defined.
@@
The relation $\leq$” on the set $\mathrm{CSC}(G)$ defined as above (for $G$ any compact Lie group) will be shown to define a partial order on $\mathrm{CSC}(G)$; hence we obtain a notion of lower semicontinuous functions to $\mathrm{CSC}(G)$. To prove antisymmetry, the following Cantor-Bernstein-like theorem is needed, which we learnt from [7].
@@
We include a proof of the theorem in order to make the article self-contained, moreover we give a slightly different argument for the case of multiple connected components compared to [7].
@@
We start out by proving:
Claim. The linear map $d_{1}f:T_{1}K\to T_{1}H$ is an isomorphism.
Proof of the claim. First, we show injectivity. It is well known that for the exponential maps $\exp_{K}$ and $\exp_{H}$ of $K$ resp. $H$ we have $\exp_{H}\circ d_{1}f=f\circ\exp_{K}$. Furthermore, both exponential maps are local diffeomorphisms, in particular injective if restricted to a suitable neighbourhood of zero in the respective tangent spaces.
@@
and therefore $w=0$ by injectivity of $(f\circ\exp_{K})|_{U}$. Hence, also $v=\lambda^{-1}w=0$ and we conclude that $d_{1}f$ is injective.
Now, analogously, one shows that $d_{1}g:T_{1}H\to T_{1}K$ is injective. By dimension counting it follows that both maps must in fact be isomorphisms, proving the claim.
Let $H_{0},K_{0}$ be the connected components of $1\in H_{0}$ and $1\in K_{0}$, respectively. It is well known that $\exp_{H}$ surjects onto $H_{0}$; moreover the claim directly states that $d_{1}f$ is surjective. Thus $f\circ\exp_{K}=\exp_{H}\circ d_{1}f$ surjects onto $H_{0}$ and since $\mathrm{im}(\exp_{K})=K_{0}$, one concludes that $H_{0}\subseteq f(K_{0})$. Hence $f|_{K_{0}}:K_{0}\to H_{0}$ is an isomorphism. The map $f$ induces a map $\overline{f}:K/K_{0}\to H/H_{0}$ which fits into the following commutative diagram with exact rows:
@@
Thus, for every $h\in H$, there exist $k_{1}\in K$ and $h_{0}\in H_{0}$ such that $h=f(k_{1})h_{0}$. On the other hand, due to surjectivity of $f|_{K_{0}}:K_{0}\to H_{0}$, there exists $k_{0}\in K_{0}$ with $h_{0}=f(k_{0})$, and then $h=f(k_{1})f(k_{0})=f(k_{1}k_{0})$ lies in the image of $f$. One concludes that $f$ is surjective, hence bijective.
Now, to finish the proof, it suffices to recall the well-known statement that a bijective Lie group homomorphism is already a Lie group isomorphism.
@@
Now, we can prove Theorem 1.3.
Proof of Theorem 1.3.
It is well known (see e. g. [6], Theorem 3.2.8) that for any $h\in\mathrm{Met}(M)$, the group $\mathrm{Hol}_{x}^{0}(h)$ is closed in $\mathrm{SO}(n)$. Hence the map from the statement of the theorem is actually well-defined. Now the statement follows from Theorem 1.1. ∎
@@
For any $k\in\mathbb{N}$, $(M,g_{k})$ is up to scaling describable as a subset of the hyperbolic space of dimension $2$. Hence, as is well known, $\mathrm{Hol}_{x}(g_{k})=\mathrm{Hol}_{x}^{0}(g_{k})\cong\mathrm{SO}(2)$ for all $x\in M$ (see e. g. [3], Proposition 10.79 and Chapter 10.K, Table 3). However $\mathrm{Hol}_{x}^{0}(g)=\{I_{2}\}\subset\mathrm{SO}(2)$ because $g$ is flat.
One may also perturb a Calabi-Yau metric without leaving the realm of Kähler metrics, that is, to obtain a sequence of metrics with holonomy $\mathrm{U}(n)$ converging to a metric with holonomy $\mathrm{SU}(n)$:
@@
Let $M$ be any $\mathrm{K}3$ surface and $\omega$ be the Kähler form of a Kähler metric on $M$. Let $\rho\in C^{\infty}(M,\Lambda^{1,1}T^{*}M)$ with $\rho]=0$ in cohomology (the existence of such a $\rho$ follows directly from the fact that $C^{\infty}(M,\Lambda^{1,1}T^{*}M)$ is infinite-dimensional while $H^{2}(M)$ is not).
By Yau’s solution to the Calabi conjecture in [10], there exists a unique Kähler metric $g_{1}$ on $M$ with Ricci form $\rho$ and such that its Kähler form is contained in the cohomology class of $\omega$. Moreover, there exists a unique Ricci-flat Kähler metric $g_{\infty}$ on $M$ with Kähler form contained in the cohomology class of $\omega$.
Let $g_{k}:=\frac{1}{k}g_{1}+\frac{k-1}{k}g_{\infty}$ for $k\in\mathbb{N}_{\geq 2}$. Then, the metrics $g_{k}$ are Kähler since convex combinations of Kähler metrics are again Kähler, and the sequence $(g_{k})_{k\in\mathbb{N}}$ converges in $\mathrm{Met}(M)$ to $g_{\infty}$. Moreover, for all $k\in\mathbb{N}$ the metric $g_{k}$ is not Ricci-flat (if it were, it would contradict the uniqueness statement from the Calabi conjecture).
Thus, the metrics $g_{k}$ all have holonomy group contained in $\mathrm{U}(2)$, but not in $\mathrm{SU}(2)$ (see e. g. [6], Proposition 4.4.2 and Proposition 6.1.1). However, $g_{\infty}$ is Ricci-flat and hence has its holonomy group contained in $\mathrm{SU}(2)\subset\mathrm{U}(2)$.
@@
On the other hand, it would be impossible to construct an example as the previous one on a closed manifold: On a compact manifold $M$ of dimension $n$, all Einstein metrics are stationary points of the Einstein-Hilbert functional $S:\mathrm{Met}(M)\to\mathbb{R}$, $g\mapsto\frac{\int_{M}\mathrm{scal}^{g}\,\mathrm{dvol}^{g}}{\left(\int_{M}\mathrm{dvol}^{g}\right)^{(n-2)/n}}$. In particular, since $S$ needs to remain constant along any deformation within the space of Einstein metrics, it is impossible to deform Ricci-flat Einstein metrics (such as hyperkähler ones) on a compact manifold to non-Ricci-flat Einstein metrics (such as quaternion-Kähler ones). Contrarily, Theorem 1.3 holds without a compactness assumption on the manifold.
To conclude this paper, it shall be remarked that one would assume that the conditions on Theorem 1.1 can actually be weakened. The most apparent restriction is the one to connections compatible with positive definite metrics, which makes the theorem, for example, inapplicable in the Lorentzian setting. Moreover, it would be interesting to see whether Theorem 1.3 also holds for the full holonomy group (which is, by [9], in general not closed in the special orthogonal group).
However, the relaxation of these conditions would require different proofs. It would be interesting to study such generalizations.
Appendix A The compact-open $C^{r}$-topology
In order for this article to be self-contained, we recall here the definition of the compact-open $C^{r}$-topology (see e. g. [5], Chapter 2.1).
Definition A.1.
Let $r\geq 0$, $M$ be a $C^{r+1}$-manifold and $F\to M$ be a vector bundle of class $C^{r}$. Given any two charts $\mathbf{x}:U\to V,\mathbf{v}:U^{\prime}\to V^{\prime}$ of $M$ respectively $F$, a compact $K\subseteq U$, $\epsilon>0$ and $f\in C^{r}(M,F)$, let
$B(f,\mathbf{x},\mathbf{v},K,\epsilon):=\Big\{$ $g\in C^{r}(M,F)\,|\,g(K)\subseteq V,\forall 0\leq k\leq r:$
$\left\lVert D^{k}(\mathbf{v}\circ f\circ\mathbf{x}^{-1})-D^{k}(\mathbf{v}\circ g\circ\mathbf{x}^{-1})\right\rVert_{C^{0}}<\epsilon\Big\}.$
The set of all such $B(f,\mathbf{x},\mathbf{v},K,\epsilon)$ has the basis property and thus defines a topology on $C^{r}(M,F)$, the compact-open $C^{r}$-topology.
If a sequence of tensors $(T_{k})_{k\in\mathbb{N}}\subset C^{0}(M,F)$ converges in the compact-open $C^{0}$-topology to a tensor $T\in C^{0}(M,F)$, then for any compact $K\subseteq M$, the restrictions $T_{k}|_{K}$ converge to $T|_{K}$ uniformly on $K$ (with respect to any set of charts covering $K$).